Project

General

Profile

11/11/2013 Minutes

Attendance

Kanika, Tricia, Michael, Susan, Nick, Dominick, Gavin, Jan

Operations

Switched to data-disk-02 for raw files at Ash River. novadaq-far cluster was full due to log files filling disk.
Three cronjobs
Michael was on shift, didn't touch FarDet at all. Smooth running. Reading everything out into Partition 1, DCMs 1,2,3 on Di-block not currently in run - power supplies died this morning.
Abbey did her threshold scans...

Production

Ask Martin how triggers are issued.

Michael: ADC spectra on top/side look normal. Definitely hotter on top than sides, geometric explanation.
Added a text box (average of points too) on plots for last entry. Added number of DCM plot.

Fernanda generated the plots requested last week. We'll discuss them tomorrow.
EventDisplay is being funky...it's masking off with not the correct mask. How?